# Online Writer Retrieval with Chinese Handwritten Phrases: A Synergistic Temporal-Frequency Representation Learning Approach

### v1

@@ III-B High Frequency Gated Attention
[…]
@@ III-C Context-Aware FPN
Subsequently, we obtain the final temporal feature embeddings $f_{T}$, classification logits $logit$, and final frequency feature embeddings $f_{F}$ through:
@@
where $concat(\cdot)$ denotes the concatenation operation, $N_{c}$ is the number of writers. $W_{l}^{T}$ is the weight of the final classification layer. $f_{T}$, $logit$, and $f_{F}$ are fed into loss computation for model optimization.
@@ III-D Model Optimization
In addition, open-set retrieval requires the model to distinguish handwriting of different writers, even though the writers are unseen during training. Therefore, we leverage the training data writers as surrogate classes to cultivate this ability. We adopt the cross-entropy loss with label smoothing $\epsilon$ as the Writer-ID loss $L_{ID}$, receiving the multi-classification logits $logit$ for optimization. This supervision allows the model to better differentiate unseen testing writers.
@@ 14 Pressure: $p$
V-B Evaluation Metric
$\centering mAP=\frac{1}{N_{q}}\sum^{N_{q}}_{i=1}AP_{i}.\@add@centering$ (21)
TABLE III: Comparison between DOLPHIN and other State-of-the-Art methods on the OLIWER dataset, the segmented CASIA-OLHWDB2 dataset, the segmented DCOH-E dataset, and the SCUT-COUCH2009 dataset. On-SV denotes online signature verification. Off-WID denotes offline writer identification. Re-ID denotes person re-identification. Off-WR denotes offline writer retrieval. On-WR denotes online writer retrieval. $\uparrow$ signifies that the higher value is better. Each experiment is repeated 50 times. Results are reported as $avg$$std$), where $avg$ is the average performance and $std$ is the standard deviation. The best performances are marked in bold and the second-best results are marked with underline.
@@ V-E Comparison with State-of-the-Art Method
We compare our proposed DOLPHIN with the State-of-the-Art (SOTA) methods on the OLIWER dataset, the segmented CASIA-OLHWDB2 dataset [13], the segmented DCOH-E dataset [14], and the SCUT-COUCH2009 dataset [15]. Due to the scarcity of established methods tailored for online writer retrieval, we draw upon SOTA methods from other domains, including the generic domain [56, 8, 60, 62, 63, 61], writer identification [16], online signature verification [17], person re-id [20, 18, 19], and offline writer retrieval [26]. Apart from Sig2Vec [17] for online signature verification, other models are originally 2D models that take images as input. To adapt these models for our task, we convert them into their 1D version by simply substituting the Conv2d module with the Conv1d module in Pytorch [66] implementations, which avoids altering their core designs and functionalities. For rigorous evaluation, we repeated each experiment 30 times, randomly selecting the query set and gallery set from the testing data as described in Sec. V-A in each iteration. The results are presented in Table III, in which we report the average $avg$ and standard deviation $std$ of the performance in the format of $avg$$std$).
@@ V-F Ablation Study
High Frequency Gated Attention (HFGA). The efficacy of the HFGA block is assessed in multiple scenarios. (1) Upon the utilization of CAIR, we integrate the Frequency Network composed of HFGA blocks into the model. This leads to a 1.6% enhancement in mAP as shown in line 3 and line 4. (2) Upon the utilization of both CAIR and C-FPN, we further add HFGA into the model and form the holistic DOLPHIN model. From line 7, this integration achieves the optimal performance, resulting in a 1.15% improvement in R1 accuracy and a substantial 3.15% improvement in mAP. (3) We investigate a variant of the HFGA block that adopts the multi-head attention mechanism [48] instead of the original single-head attention mechanism, denoted as HFGA-MH. The results are presented in Table V. Interestingly, adopting multi-head attention leads to performance degradation, with the added drawback of consuming 3×3$\times$3 × GPU memory for training the same architecture model. Therefore, we utilize single-head attention for better performance and computational efficiency. The above experiments underscore the effectiveness of the proposed HFGA block, which could be attributed to HFGA’s synergy between temporal and frequency features via the gated cross-attention mechanism. By attending to the informative high-frequency components and facilitating interaction between the temporal and frequency domains, HFGA distills a robust understanding of the writer’s unique stylistic traits, thus enhancing the model’s performance.

### v2

@@ III-B High Frequency Gated Attention
Frequency analysis excels at revealing discriminative writing features that usually remain hidden in the time domain, such as rhythmic patterns and scale-invariant stroke formations that characterize an individual’s unique handwriting style. Nevertheless, frequency features have been largely overlooked in online handwriting modeling. To this end, we propose the High Frequency Gated Attention (HFGA) block to capitalize on the rich and distinctive writing features in the high-frequency spectrum. Fig. 4 depicts the schematic of the HFGA block. Formally, considered a handwritten sample in the temporal domain $x^{t}\in\mathbb{R}^{L\times d}$, we utilize the 1D Discrete Wavelet Transform (DWT) to decompose $x^{t}$ into the high frequency components $x^{f}_{h}\in\mathbb{R}^{{\lfloor L/2\rfloor}\times d}$ and low frequency components $x^{f}_{l}\in\mathbb{R}^{{\lfloor L/2\rfloor}\times d}$. Here, $x^{f}_{h}$ corresponds to the sharp turns and fine-grained details of the signal, which represent local features such as stroke curvatures, pressure variations, and writing speed fluctuations. $\lfloor\cdot\rfloor$ denotes the flooring operation.
Figure 5: Schematic of the Context-Aware FPN.
@@ III-C Context-Aware FPN
Subsequently, we obtain the final temporal feature embeddings $f_{T}$, classification logits logitlogit, and final frequency feature embeddings $f_{F}$ through:
@@
where $concat(\cdot)$ denotes the concatenation operation, $N_{c}$ is the number of writers. $W_{l}^{T}$ is the weight of the final classification layer. $f_{T}$, logitlogit, and $f_{F}$ are fed into loss computation for model optimization.
@@ III-D Model Optimization
In addition, open-set retrieval requires the model to distinguish handwriting of different writers, even though the writers are unseen during training. Therefore, we leverage the training data writers as surrogate classes to cultivate this ability. We adopt the cross-entropy loss with label smoothing $\epsilon$ as the Writer-ID loss $L_{ID}$, receiving the multi-classification logits logitlogit for optimization. This supervision allows the model to better differentiate unseen testing writers.
@@ IV OLIWER Dataset
Algorithm 1 Procedure for segmenting online handwritten lines into online handwritten phrases, with only stroke-level annotations of the data.
  • (1)
[…]
  • (2)
    Since timestamps are pre-known, we can calculate the time interval $ts_{i-1}$ between stroke $s_{i-1}$ and stroke $s_{i}$ by subtracting the first timestamp of $s_{i}$ from the last timestamp of $s_{i-1}$, and get a sequence of time intervals $T=\{ts_{1},ts_{2},...,ts_{N-1}\}$.
  • (3)
    Given the text transcription $Y$ of line $X$, we calculate the number of Chinese characters in the transcription, denoting as $N_{Y}$. We sort $T$ in descending order and select the $ts$ ranked in $N_{Y}+1$ position as a time threshold $d$. In this case, there should be $N_{Y}$ time intervals $\textgreater~d$ and we denote them as $T^{d}=\{ts_{1}^{d},ts_{2}^{d},...,ts_{N_{Y}}^{d}\}$, while the rest are $\leq d$.
  • (4)
    We sequentially traverse all the strokes in $X$ and utilize the time intervals in $T^{d}$ as division points to cluster the strokes between two division points, in which each cluster represents an isolated character. This leads to in total $N_{Y}$ isolated characters segmented from the given text line $X$.
@@ 14 Pressure: $p$
TABLE III: Comparison between DOLPHIN and other State-of-the-Art methods on the OLIWER dataset, the segmented CASIA-OLHWDB2 dataset, the segmented DCOH-E dataset, and the SCUT-COUCH2009 dataset. On-SV denotes online signature verification. Off-WID denotes offline writer identification. Re-ID denotes person re-identification. Off-WR denotes offline writer retrieval. On-WR denotes online writer retrieval. $\uparrow$ signifies that the higher value is better. Each experiment is repeated 50 times. Results are reported as avgavgstdstd), where avgavg is the average performance and stdstd is the standard deviation. The best performances are marked in bold and the second-best results are marked with underline.
@@ V-E Comparison with State-of-the-Art Method
We compare our proposed DOLPHIN with the State-of-the-Art (SOTA) methods on the OLIWER dataset, the segmented CASIA-OLHWDB2 dataset [13], the segmented DCOH-E dataset [14], and the SCUT-COUCH2009 dataset [15]. Due to the scarcity of established methods tailored for online writer retrieval, we draw upon SOTA methods from other domains, including the generic domain [56, 8, 60, 62, 63, 61], writer identification [16], online signature verification [17], person re-id [20, 18, 19], and offline writer retrieval [26]. Apart from Sig2Vec [17] for online signature verification, other models are originally 2D models that take images as input. To adapt these models for our task, we convert them into their 1D version by simply substituting the Conv2d module with the Conv1d module in Pytorch [66] implementations, which avoids altering their core designs and functionalities. For rigorous evaluation, we repeated each experiment 30 times, randomly selecting the query set and gallery set from the testing data as described in Sec. V-A in each iteration. The results are presented in Table III, in which we report the average avgavg and standard deviation stdstd of the performance in the format of avgavgstdstd).
@@ V-F Ablation Study
High Frequency Gated Attention (HFGA). The efficacy of the HFGA block is assessed in multiple scenarios. (1) Upon the utilization of CAIR, we integrate the Frequency Network composed of HFGA blocks into the model. This leads to a 1.6% enhancement in mAP as shown in line 3 and line 4. (2) Upon the utilization of both CAIR and C-FPN, we further add HFGA into the model and form the holistic DOLPHIN model. From line 7, this integration achieves the optimal performance, resulting in a 1.15% improvement in R1 accuracy and a substantial 3.15% improvement in mAP. (3) We investigate a variant of the HFGA block that adopts the multi-head attention mechanism [48] instead of the original single-head attention mechanism, denoted as HFGA-MH. The results are presented in Table V. Interestingly, adopting multi-head attention leads to performance degradation, with the added drawback of consuming $3\times$ GPU memory for training the same architecture model. Therefore, we utilize single-head attention for better performance and computational efficiency. The above experiments underscore the effectiveness of the proposed HFGA block, which could be attributed to HFGA’s synergy between temporal and frequency features via the gated cross-attention mechanism. By attending to the informative high-frequency components and facilitating interaction between the temporal and frequency domains, HFGA distills a robust understanding of the writer’s unique stylistic traits, thus enhancing the model’s performance.